Circular No. 5765

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA OPHIUCHI 1993 Paul J. Camilleri, Cobram, Vic., Australia, reports his discovery of an

apparent nova on T-Max patrol films taken with an 85-mm camera lens. The object is located at α = 17^{h}22^{m}.1, δ = −23^{o}09′ (equinox 1950.0). The following magnitude estimates are available (photovisual unless otherwise noted): Apr. 1.77 UT, [12 (Camilleri); 14.54, 9.5 (Camilleri); 16.68, 9.5 (Camilleri); 17.50, 9.5 (P. F. Williams, Heathcote, N.S.W.; visual); 17.52, 9.5 (Camilleri; visual). There is no image visible at this location on Atlas Stellarum (1970 Apr. 6) or the Union Observatory Prints (1921 June 4) to a limiting mag of 14. The new object is ∼ 2′ east of a star of mag 10.

SUPERNOVA 1993J IN NGC 3031 Further photoelectric photometry (cf. IAUC 5750, 5755) by D. Hanzl,

N. Copernicus Planetarium, Brno (H); by B. Kulesza, Olsztyn Observatory, Poland (K); by M. Mikołajewski, K. Rumiński, and M. Staniewski, N. Copernicus University, Toruń (M); and by P. Prugniel, Observatoire de Haute Provence (P): Apr. 13.84 UT, V = 11.07 ± 0.03, B–V = +0.56 ± 0.06, V–R = +0.50 ± 0.05, V–I = +0.42 ± 0.06 (M); 14.00, B_{c} = 11.70, B_{c}–R_{c} = +0.87 (P); 14.89, V = 10.88 ± 0.04, B–V = +0.56 ± 0.07 (H); 14.91, V = 10.82 ± 0.04, B–V = +0.55 ± 0.03 (K); 14.93, V = 10.94 ± 0.01, U–B = –0.06 ± 0.04, B–V = +0.51 ± 0.02, V–R = +0.48 ± 0.02, V–I = +0.53 ± 0.02 (M); 15.08, B_{c} = 11.65, B_{c}–R_{c} = +0.89 (P); 16.00, B_{c} = 11.62, B_{c}–R_{c} = +0.89 (P); 17.00, B_{c} = 11.63, B_{c}–R_{c} = +0.93 (P). Using HD 86677 (cf. IAUC 5742), Hanzl finds V = 11.24 ± 0.04, B–V = +0.67 ± 0.03 for GSC 4383.0565.

BL LACERTAE The National Astronomical Observatory in Tokyo has forwarded the

following report from M. Okyudo, Nishi-Harima Astronomical Observatory: “BL Lac has been highly active since the latter half of 1992, when its magnitude was V ∼ 15. Recent CCD photometric observations (Johnson system): 1993 Apr. 5.77 UT, V = 14.58 ± 0.05; 14.77, 13.78 ± 0.02. Comparison stars are ‘B’ and ‘C’ of Smith et al. (1985, A.J. 90, 1184), which are in the same field as BL Lac.”

1993 April 17 Daniel W. E. Green